270g+5(15-g)=164(15)
Found 2 solutions by tutorcecilia, acerX:
Answer by tutorcecilia(2152)

270g+5(15-g)=164(15)
270g+75-5g=2460
270g+75-75-5g=2460-75
270g-5g=2385
265g=2385
g=2385/265
g=477/53
g=9
.
Check by plugging (g=9) back into the original equation.

Answer by acerX(62)

270g+5(15-g)=164(15)
Distribute:
270g+75-5g=2460
Isolate variables:
270g-5g=2460+75
Simplify:
265g=2535
Divide both sides by 265:
g=35